RC meeting and Run Plan, Sunday December 4th
From SBS wiki
Return to the RC Meeting Calendar
Contents
RC Meeting Information
Data taking has been smooth. No RC meeting for today, Have a nice weekend!
Announcement
Bill will host a target training session on next Monday after the RC meeting. We have lots of TO shifts open; please take the chance to join the session
Current Status
- Very smooth production run on He3 target @ 36 uA. No big interruptions.
- NMR polarization - 46.8% (rough) at the time of this log entry
- During Saturday Swing shift, MCC temporarily dropped us to 30 uA due to RF instability issue, I checked with MCC, and confirmed that this is only temporary, as long as they have better RF stability, they will bring us back to 36 uA.
- Jiwan, Alexander, Bryan and Ben traceing the cause of the HCal f1TDC 1 & 4 timing shift issue (see a log https://logbooks.jlab.org/entry/4095193 ). Up to now, this issue only shows up in a single run 3005. Jiwan put specific instructions on what to do when we encounter this issue again. We will keep an eye on it.
- There are a few runs with 50 K replay missing SBS GEMs plots due to the oversized pdf file, which blocked them from being posted to the logbook. Sean reduced the pdf file size by removing a few histograms. The later runs seem fine since then.
- For future shift takers, MCC gave us an option for checking beam restoration status. Please check a detailed description in a log by Iuliia [ https://logbooks.jlab.org/entry/4094990 https://logbooks.jlab.org/entry/4095153 ]
- Target increased laser power from 144 W to 155 W, please see a summary log entry by Jack ( https://logbooks.jlab.org/entry/4095195 )
- Paul ran the charge asymmetry feedback briefly during run 10038, please see log entries by Paul ( https://logbooks.jlab.org/entry/4095262 https://logbooks.jlab.org/entry/4095247 https://logbooks.jlab.org/entry/4094291 )
- No major issue in OWL shift
- Sunday morning, MCC brought us back to 36 uA beam
Short Term Run Plan
- Continue production run on He3 target, 36 uA
- Use GEnII-3stream DAQ config
- Use ps3=0, ps5=6, others =-1
- BBCal threshold at -366 mV, HCal threshold at -19 mV
- NMR measurement every 2 hours
Current Issues
- BB Hodoscope: channel 0R - 15R missing in the 50K replay [ https://logbooks.jlab.org/files/2022/12/4093673/summaryPlots_2966_BBTH.pdf ]
- One APV malfunction, not a big problem
Planned Accesses to the Hall
- Will try to arrange an escorted access for fixing hodoscope when we have a long beam down ( longer than 1 hour).
Special Requests
- None
Notices in Effect for Shift Crew
- Power down SBS, BB, and Detector HV before making an access to the Hall